### v1

# FinReflectKG - EvalBench: Benchmarking Financial KG with Multi-Dimensional Evaluation

Abstract
Large language models (LLMs) are increasingly being used to extract structured knowledge from unstructured financial text. Although prior studies have explored various extraction methods, there is no universal benchmark or unified evaluation framework for the construction of financial knowledge graphs (KG). We introduce FinReflectKG - EvalBench, a benchmark and evaluation framework for KG extraction from SEC 10-K filings. Building on the agentic and holistic evaluation principles of FinReflectKG - a financial KG linking audited triples to source chunks from S&P 100 filings and supporting single-pass, multi-pass, and reflection-agent-based extraction modes - EvalBench implements a deterministic commit-then-justify judging protocol with explicit bias controls, mitigating position effects, leniency, verbosity and world-knowledge reliance. Each candidate triple is evaluated with binary judgments of faithfulness, precision, and relevance, while comprehensiveness is assessed on a three-level ordinal scale (good, partial, bad) at the chunk level. Our findings suggest that, when equipped with explicit bias controls, LLM-as-Judge protocols provide a reliable and cost-efficient alternative to human annotation, while also enabling structured error analysis. Reflection-based extraction emerges as the superior approach, achieving best performance in comprehensiveness, precision, and relevance, while single-pass extraction maintains the highest faithfulness. By aggregating these complementary dimensions, FinReflectKG - EvalBench enables fine-grained benchmarking and bias-aware evaluation, advancing transparency and governance in financial AI applications.
Keywords: Knowledge Graphs, LLM-as-a-Judge, Evaluation Benchmarks, SEC Filings, Large Language Models, Natural Language Processing, Financial AI
1 Background and Motivation
Large language models (LLMs) are increasingly used in finance across diverse tasks, including the extraction of information from long and unstructured documents. A common target representation for such extracted knowledge is the KnowledgeGraph (KG), which organizes information into structured triples and supports downstream applications such as compliance monitoring, risk management, and large-scale financial analytics [1, 2]. Although recent advances in LLM have improved triple extraction through multi-turn prompting and reflection [3], the evaluation of extracted triples remains a critical challenge [4]. Since financial KGs are designed to support high-stakes applications such as investment research, credit risk assessment, and portfolio decisions, the lack of reliable validation for LLM-generated triples poses significant risks.
Previous works have explored LLM-as-Judge as a scalable alternative to costly human annotation [5, 6]. However, LLM judges are susceptible to prompt sensitivity and order-based position effects [7], often favoring answers presented earlier, regardless of quality. They also tend to be overly lenient [8] and exhibit stylistic biases that reward persuasive phrasing over factual correctness [9], which undermines the robustness and interpretability of LLM-based evaluations [10, 11]. Recent studies on prompt design show that instruction-only prompting reduces reasoning drift and improves judgment consistency [12, 13, 14], whereas encouraging detailed chain-of-thought often produces the opposite effect, leading to overthinking and topic drift [15]. Moreover, evidence suggests that few-shot prompting further stabilizes LLM-as-Judge behavior and increases agreement with human annotations [16]. . Although several open-source toolkits (e.g., DeepEval, Ragas, TruLens, LangSmith) incorporate LLM-as-Judge components for tasks such as QA and summarization [17, 18, 19, 20], rigorous, bias-aware evaluation protocols tailored to triple-level KG extraction, particularly in finance, remain underspecified.
To address these challenges, we introduce FinReflectKG - EvalBench, a benchmark for financial KG extraction built on SEC 10-K filings. FinReflectKG - EvalBench extends the agentic principles of FinReflectKG [21] by combining schema-aware extraction with a conservative evaluation protocol. To the best of our knowledge, this is the first systematic benchmark for financial KG extraction that integrates various extraction strategies (single pass, multi pass and reflection) and a rigorous evaluation framework.
This paper makes the following key contributions:
  1. 1.
    We present FinReflectKG - EvalBench, the first benchmark for financial KG extraction from SEC filings with a reproducible and bias-aware evaluation framework.
  2. 2.
    We provide a comparative evaluation of extraction modes (single-pass, multi-pass, and reflection) across four complementary dimensions: faithfulness, precision, relevance, and comprehensiveness.
2 Methodology
We construct FinReflectKG - EvalBench on the corpus of U.S. SEC Form 10-K filings from S&P 100 companies for fiscal year 2024. Formally, let $\mathcal{D}$ denote this corpus where each document $d\in\mathcal{D}$ is segmented into text spans $\mathcal{X}_{d}=\{x_{d,1},\dots,x_{d,K_{d}}\}$ using a deterministic, structure-aware chunking scheme tailored to financial filings. An information extractor $E$ maps each span to a set of candidate triples,
$T_{x}=E(x)\subseteq\mathcal{T},\qquad t=(s,r,o)\in\mathcal{T},$
where $s$ and $o$ are subject and object entity mentions, and $r\in\mathcal{R}$ is a relation from a predefined financial vocabulary.
We evaluate extraction outputs using an LLM-as-Judge $J$, instantiated with the Qwen3-32B model and configured for deterministic decoding (temperature =0.0=0.0). As shown in Figure 1, and following the commit-then-justify paradigm [22], the judge first produces a structured verdict and then a concise justification (up to 15 words). To further support error analysis, we introduce a warning signal that highlights extraction errors and provides actionable correction paths. These signals can also be leveraged within a feedback loop for iterative self-improvement.
To ensure reliable and reproducible judgments, we enforce strict bias controls. First, we adopt a principle of conservatism: whenever the evidence is ambiguous, the judge defaults to a negative decision (0), thereby mitigating leniency bias. Second, we enforce locality, strictly prohibiting the use of world knowledge or inferences beyond the provided text. Third, we guarantee position independence by instructing the judge not to let the order or placement of sentences influence its verdicts. Finally, we ensure verbosity independence, so that the length or surface form of a candidate triple does not bias the evaluation outcome. In addition, we improve consistency and calibration by including few-shot examples for each evaluation criterion in the judge’s prompt. Concrete examples are provided in the Appendix.
role: "Knowledge Graph Evaluator"
task: "Determine if the triplet is [evaluation criterion]
with respect to the source text context"
instructions:
Decision rule:
- Return 1 if [criterion satisfied],
- Return 0 if [criterion not satisfied].
Bias controls:
- Be conservative: when uncertain, return 0 (leniency bias).
- Do NOT infer or add information beyond the text (world knowledge bias).
- Do NOT let sentence position in the source text affect the decision (position bias).
- Do NOT let the length of the triplets affect the decision (verbosity bias).
Reasoning vs Warning:
- Reasoning: Brief explanation of the verdict (up to 15 words).
- Warning: actionable tag(s) for error type; Do NOT duplicate reasoning; use empty string if no actionable issue.
Output policy:
- Valid JSON array only, single line.
- Each item: {"verdict":0|1,"reasoning":"...","warning":"..."}
Examples:
Figure 1: System prompt design
The three extraction modes define different approaches to constructing knowledge graph triples. Single-pass uses a single LLM for both extraction and normalization. Multi-pass splits the task between two LLMs: one extracts triples, while the other normalizes them according to rubric parameters. Reflection employs an agentic, iterative workflow, where extraction and feedback loops refine triples until inconsistencies are resolved or a maximum iteration limit is reached. Across the three extraction modes, the task is to evaluate the quality of the candidate triples produced for each span $x$ with candidate set $T_{x}$. Evaluation is conducted along four complementary dimensions. First, faithfulness $F$ measures whether the content of a triple is factually grounded in the source text, without relying on world knowledge or bridging inferences. Second, precision $P$ assesses the clarity and specificity of triples, penalizing generic placeholders (e.g., “Company”) and imprecise expressions of quantities or dates. Third, relevance $R$ checks whether the triple contributes directly to the main theme of the source span rather than introducing tangential information. Finally, comprehensiveness $C$ is measured on a three-level ordinal scale (good, partial, bad) and it evaluates coverage at the chunk level, scoring how well the set of triples represents all atomic core facts.
To aggregate results across the corpus, let $\mathcal{X}=\bigcup_{d\in\mathcal{D}}\mathcal{X}_{d}$ and $\mathcal{T}_{\text{all}}=\bigcup_{x\in\mathcal{X}}T_{x}$. Local binary metrics (faithfulness, precision, relevance) are micro-averaged:
$\bar{F}=\frac{1}{|\mathcal{T}_{\text{all}}|}\sum_{(x,t)}F(x,t),\qquad\bar{P},\;\bar{R}\text{ analogously}.$
Comprehensiveness is macro-averaged across spans:
$\bar{C}=\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}C(x).$
3 Results
Evaluation of triple generation often treats faithfulness, relevance, precision, and comprehensiveness as separate dimensions. While such a decomposition is informative, we argue that these criteria must ultimately be interpreted jointly in order to capture the full spectrum of trade-offs involved in knowledge graph construction.
Table 1: Evaluation results across extraction modes
Single Pass Multi Pass Reflection
Comprehensiveness 62.60 62.41 72.01
Faithfulness 87.25 78.73 83.40
Precision 56.06 58.01 59.49
Relevance 91.46 82.64 92.52
As reported in Table 1, performance varies markedly across extraction modes. The reflection mode achieves superior results in comprehensiveness, precision, and relevance, whereas the single-pass mode yields the highest score in faithfulness. This pattern is consistent with the intuition that reflection, by design, generates a larger set of triples per chunk, thereby capturing a broader range of atomic core facts. From the perspective of knowledge graph construction, this ability to recover a wider coverage of facts is highly desirable, as it directly impacts the downstream utility of the graph.
At the same time, the relative decline in faithfulness observed for reflection suggests an inherent trade-off: expanding coverage increases the risk of generating triples that extend beyond the strict boundaries of the source text. In contrast, the single-pass approach, though less comprehensive, remains more conservative and better aligned with the original text.
Turning to precision, we observe that absolute scores remain relatively modest across all modes, indicating the need for further improvements. Nonetheless, reflection achieves the best performance in this dimension, suggesting that iterative reasoning can modestly improve structural accuracy. By contrast, relevance yields consistently higher values across modes, with reflection leading, which indicates that most generated triples-despite occasional issues of faithfulness or precision-remain topically aligned with the source text.
Together, these results highlight the complementary strengths and weaknesses of different extraction strategies. Reflection provides the most balanced coverage, excelling in breadth, topical alignment, and structural accuracy, though it requires further calibration to match the strict factual reliability of single-pass generation. These findings underscore the importance of multi-dimensional evaluation, as no single metric alone captures the full spectrum of trade-offs inherent to financial KG construction.
4 Conclusions
In this work, we introduced FinReflectKG - EvalBench, the first benchmark for financial knowledge graph extraction from SEC 10-K filings with a reproducible, bias-aware evaluation framework. By integrating schema-aware extraction with a conservative LLM-as-Judge protocol, we provide a principled way to assess triple quality across multiple extraction modes. Our analysis highlights several key insights. Reflection emerges as the most balanced modes, achieving superior comprehensiveness, precision, and relevance, while single-pass remains the most faithful to the source text. This trade-off underscores the necessity of multi-dimensional evaluation, as no single metric alone can fully capture extraction quality.
Beyond reporting scores, FinReflectKG - EvalBench contributes a methodological advancement through explicit bias controls, a commit-then-justify judgment protocol, and the introduction of warning signals that enable actionable error analysis. These design elements not only improve reliability but also open the door to self-improving extraction pipelines, where diagnostic signals guide iterative refinement.
Looking forward, our benchmark provides a foundation for advancing financial KG research in several directions. Future work may extend coverage beyond the S&P 100 universe and incorporate diverse categories of financial documents. By establishing transparent, reproducible, and bias-aware evaluation standards, FinReflectKG - EvalBench aims to accelerate progress toward trustworthy financial KGs that support downstream tasks in compliance, risk management, and large-scale financial analytics.